OpenStudy (anonymous):

What's the answer to -6 exponet 0 times 1 over 3 negative exponet 2

OpenStudy (smored):

\[-6^{0}*\frac{ 1 }{ 3 }^{-2}\] is this it?

OpenStudy (anonymous):

exponet on the bottom

OpenStudy (anonymous):

\[\Large-6^0*\frac{1}{3^{-2}}\]So like this?

OpenStudy (smored):

\[-6^{0}=-(6^{0})=-1\] any non-zero number with exponent 0 = 1 \[\frac{ 1 }{ 3^{-2} } = \frac{ 3 }{ 1 }^{2}=9\] a number with negative exponent is its reciprocal with the positive exponent -1*9=-9

OpenStudy (anonymous):

yes, thanks

OpenStudy (smored):

no prob